if the average of 5 consecutive odd numbers is 11, then the largest number is what

Answer by unlockmath(1688)

Hello,
Let x represent the first odd number. Then the other 4 will be shown as:
[x+(x+2)+(x+4)+(x+6)+(x+8)]/5=11
Combine like terms to get:
(5x+20)/5=11
Multiply both sides by 5 to get:
5x+20=55
Subtract 20 to get:
5x=35
Divide by 5 to get:
x=7
So the largest is 15
Make sense?
RJ
www.math-unlock.com